### v1

@@ Abstract
Dexterous manipulation, Sim2Real, articulated object, world model
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I INTRODUCTION
Articulated object manipulation is a fundamental and challenging problem in robotics. Compared with pick-and-place tasks, where only the start and final poses of robot end effectors are constrained, articulated object manipulation requires the robot end effector to move along certain trajectories, making the problem significantly more complex. Most existing works utilize a neural network to learn the correlation between object states and correct actions, and employ reinforcement learning (RL) and imitation learning (IL) to train the neural network [1, 2, 3]. However, since the state distribution of articulated objects is higher-dimensional and more complex than that of rigid objects, it is difficult for the neural network to learn such correlation, even with hundreds of successful demonstrations and millions of interactions [4, 5].
@@
For humans, manipulation involves not only action responding to perception, as is the case with policy networks, but also motor imagery and mental simulation, that humans can imagine the action consequences before execution and plan the action trajectory accordingly[6]. To model the world more accurately, humans can actively interacting with the environment, changing its states and gathering additional information, which is named as interactive perception [7, 8].
@@
While dexterous manipulation with multi-finger hands enables more flexible, efficient and robust manipulation, the high-dimensional action space presents significant challenges for MPC. To handle this problem, we propose to employ eigengrasp [9] to reduce the operational dimensions of the dexterous hand, enabling more efficient and successful searching. While eigengrasp has been widely studied for robot grasping [10, 11, 12], its application in dexterous manipulation remains under-explored. Since our method constructs an explicit world model of the articulated object, we can accurately predict its motion upon contact with the dexterous hand. This allows us to search for a feasible dexterous manipulation trajectory.
@@
(1) We broaden the framework’s scope from manipulation with two-finger gripper to multi-finger dexterous manipulation. To address the challenge introduced by the high-dimensional action space of the dexterous hand, we propose to utilize eigengrasp to reduce the dimension, leading to more efficient and successful manipulation. We conduct extensive experiments both in simulation and on a real robot to validate our method’s effectiveness for dexterous manipulation and the usefulness of its different modules.
@@
The remainder of this paper is structured as follows: Related works are reviewed in Section II. Our proposed robot learning framework is detailed in Section III. Experimental setup and results are presented in Section IV. Finally, conclusions, limitations, and future directions are discussed in Section V.
@@ II RELATED WORK
II-A Dexterous Manipulation
@@
II-B World Model Construction
@@
In our work, we focus on articulated object manipulation, so we introduce the knowledge prior of the environment by using an explicit physics model. Therefore, we are able to decrease the number of samples required for model construction to 1. Moreover, the explicit physics model’s generalizability guarantees that while we only use a simple action to collect the sample, the built model can be used for long-horizon complex trajectory planning composed of unseen robot actions.
II-C Affordance Learning
@@
II-D Sim2Real for Robot Learning
@@ III METHOD
[…]
The goal of our work is to manipulate articulated objects to specified joint states with various robot-effectors in the real world, including two-finger grippers and multi-finger dexterous hands. To better align with actual application scenarios, we employ a single depth sensor to acquire a partial point cloud of the object as the observation. Fig. 2 shows an overview of our framework. It consists of three modules: Interactive Perception (Section III-A), Explicit World Model Construction (Section III-B), Sampling-based Model Predictive Control (Section III-C).
A single observation of an articulated object cannot provide enough information to reveal its full structure. For example, when humans first look at a kitchen door, it is hard to tell whether it has a rotating hinge or a sliding hinge. However, after the door is moved, humans can use the information from the two observations to infer the type and location of the hinge. Inspired by this, the Interactive Perception module proposes an action to alter the joint state of the articulated object based on learned affordance. This action is then executed on the object in the real world, resulting in two frames of point clouds: one before the interaction and one after.
With the two point clouds, the Explicit World Model Construction module (Section III-B) infers the shape and the kinematic structure of the articulated object to construct a digital model. The digital model can be loaded into a physics simulator for the robot to interact with, forming an explicit world model of the environment.
@@ III-A2 Learn from real-world egocentric demonstrations
Acquiring 3D affordance representations through self-supervised interactions in simulation has shown promise as it doesn’t rely on labeled data. However, certain limitation exists: the success of this method hinges on interactive models in simulation. Unfortunately, the availability of simulated datasets for articulated objects is limited, hindering the generation of training data.
@@
Fig. 3 illustrates how we generate a 3D trajectory for real robot manipulation from 2D affordances. Firstly, we capture an RGB image $\boldsymbol{I}_{0}$ and a 3D point cloud $\bm{\mathcal{P}}_{0}$ using the mounted RGBD camera, $\boldsymbol{H}_{0}\in\mathbb{R}^{4\times 4}$ is its relative transformation matrix respect to the robot coordinate system. Secondly, we set a virtual camera, the relative transformation matrix of which is $\boldsymbol{H}\in\mathbb{R}^{4\times 4}$. Since the depth of each pixel in $\boldsymbol{I}_{0}$ is known, we can generate the virtual RGB image $\boldsymbol{I}_{1}$ by image wrapping. Thirdly, we use $\boldsymbol{I}_{0}$ and $\boldsymbol{I}_{1}$ as the input of the affordance model and generate contact points $\boldsymbol{c}_{0}=(u_{0},v_{0})$, $\boldsymbol{c}_{1}=(u_{1},v_{1})$ and post-contact trajectories $\boldsymbol{\tau}_{0}=(u_{0}^{{}^{\prime}}-u_{0},v_{0}^{{}^{\prime}}-v_{0})$ and $\boldsymbol{\tau}_{1}=(u_{1}^{{}^{\prime}}-u_{1},v_{1}^{{}^{\prime}}-v_{1})$. The camera intrinsic matrix is $\boldsymbol{K}$, the contact point in the mounted camera frame is $\boldsymbol{p}_{c}\in\mathbb{R}^{3}$, the 3D post contact vector in the camera frame is $\boldsymbol{\tau}_{c}\in\mathbb{R}^{3}$. Fourthly, we respectively calculate the 3D contact point and post contact vector. We use contact point $\boldsymbol{c}_{0}$ to acquire 3D contact point $\boldsymbol{p}_{c}\in\mathbb{R}^{3}$ in the robot base frame:
@@
where $z_{c}$ represents the depth of $\boldsymbol{p}_{c}$. We use camera’s intrinsic matrix to transfer $\boldsymbol{c}_{0}$ to point in mounted camera frame than use mounted camera’s extrinsic matrix to transfer it to the 3D point cloud $\boldsymbol{p}_{c}$.
@@
Specifically, our method of calculating 3D post contact vector is shown in Fig. 3. We respectively denote the projection plane of $\boldsymbol{I}_{0}$ and $\boldsymbol{I}_{1}$ as $\boldsymbol{S}_{0}$ and $\boldsymbol{S}_{1}$. For $\boldsymbol{S}_{0}$, we use $\boldsymbol{\varphi}_{0}$ and $\boldsymbol{\varphi}_{0}^{{}^{\prime}}$ to represent the projection plane. $\boldsymbol{\varphi}_{0}$ represents one possible 3D vector on projection plane $\boldsymbol{S}_{0}$. Its starting point is $\boldsymbol{p}_{c}$, while its ending point can be calculated with:
@@
where $\boldsymbol{o}_{c0}$ is the coordinate of camera frame’s origin in the robot base frame. Then we calculate the norm vector of $\boldsymbol{S}_{0}$: $\boldsymbol{n}_{0}$ = $\boldsymbol{\varphi}_{0}\times\boldsymbol{\varphi}_{0}^{{}^{\prime}}$. We can calculate $\boldsymbol{n}_{1}$ in the same way. Finally, we generate the 3D post-contact vector in the robot base frame: $\boldsymbol{\tau}_{c}$ = $\boldsymbol{n}_{0}\times\boldsymbol{n}_{1}$.
Finally, we use motion planning to conduct the one-step interaction with the articulated object. The motion planning process can be divided into two phases: we first let the hand move to the contact point and then we let the hand move a little distance in the direction of the post contact vector.
@@ III-B Explicit World Model Construction
Building an explicit model of an articulated object is difficult because only if the geometries of all parts and kinematic relationships between connected parts are both figured out can the model of the articulated object be constructed.
In our work, we have two assumptions for the articulated objects: (1) the articulated object only contains a single prismatic or revolute joint; (2) the base link of the articulated object is fixed.
We choose Ditto [49] to construct the physical model explicitly. Given the visual observations before and after the interaction ($\bm{\mathcal{P}}_{0}$ and $\bm{\mathcal{P}}_{1}$), Ditto uses structured feature grids and unified implicit neural representation to construct part-level digital twins of articulated objects. Different from the original work where a multi-view fused point cloud is used, we use a single-view point cloud as input, which is more consistent with real robot application settings. Furthermore, we simulate the depth sensor’s noise when generating training data to narrow the domain gap [50]. After we train the Ditto on simulated data, we use the trained model on the real two-frame point clouds to generate the implicit neural representation and extract the meshes. The explicit physics model is represented as the Unified Robot Description Format (URDF), which can be easily loaded into widely used multi-body physics simulators, such as SAPIEN [51].

### v2

@@ Abstract
Dexterous manipulation, Sim2Real, articulated object, world model, 3D AIGC
@@ I INTRODUCTION
Articulated object manipulation is a fundamental and challenging problem in robotics. Compared with pick-and-place tasks, where only the start and final poses of robot end effectors are constrained, articulated object manipulation requires the robot end effector to move along certain trajectories, making the problem significantly more complex. Most existing works utilize a neural network to learn the correlation between object states and correct actions, and employ reinforcement learning (RL) and imitation learning (IL) to train the neural network [1, 2, 3]. However, since the state distribution of articulated objects is higher-dimensional and more complex than that of rigid objects, it is difficult for the neural network to learn such correlation, even with hundreds of successful demonstrations and millions of interactions [4, 5, 6].
@@
For humans, manipulation involves not only action responding to perception, as is the case with policy networks, but also motor imagery and mental simulation, that humans can imagine the action consequences before execution and plan the action trajectory accordingly[7]. To model the world more accurately, humans can actively interact with the environment, change its state and gather additional information, which is named as interactive perception [8, 9].
@@
While dexterous manipulation with multi-finger hands enables more flexible, efficient and robust manipulation, the high-dimensional action space presents significant challenges for MPC. To handle this problem, we propose to employ eigengrasp [10] to reduce the operational dimensions of the dexterous hand, enabling more efficient and successful searching. While eigengrasp has been widely studied for robot grasping [11, 12, 13], its application in dexterous manipulation remains under-explored. By leveraging eigengrasp and the constructed explicit world model, we can accurately predict the motion of the articulated object upon contact with the dexterous hand and search for a feasible dexterous manipulation trajectory.
@@
(1) We propose a novel articulated object model construction pipeline. Compared with Ditto [15], our pipeline is able to handle objects with more than one movable parts. Moreover, since we utilize 3D AIGC to reconstruct the geometry, the reconstruction quality is significantly improved.
(2) We broaden the framework’s scope from manipulation with two-finger gripper to more types of end effectors, including suction gripper and multi-finger dexterous hand. To address the challenge introduced by the high-dimensional action space of the dexterous hand, we propose to utilize eigengrasp to reduce the dimension, leading to more efficient and successful manipulation. We conduct extensive experiments both in simulation and on a real robot to validate our method’s effectiveness for dexterous manipulation and the usefulness of its different modules.
@@
The remainder of this paper is structured as follows: Related works are reviewed in Section II. Our proposed robot learning framework is detailed in Section III. Experimental setup and results are presented in Section IV. Experiment analysis and limitations are discussed in Section V. Finally, our work and future directions are concluded in Section VI.
@@ II RELATED WORK
II-A Learning-Based Articulated Object Manipulation
Articulated object manipulation remains an open and challenging research problem due to the complex kinematic constraint. With the development to 3D learning algorithms, larger-scale articulated object dataset and robot physics simulation platform, many learning-based works have been proposed for affordance learning [16, 17], articulation parameter estimation [18, 19, 20], and manipulation skill learning [5, 21, 22, 23, 2].
In FlowBot [21], the per-point 3D articulation flow of the point cloud is first estimated, then the point with the maximum flow vector magnitude is chosen for manipulation with a suction gripper. FlowBot++ [23] enhances this process by explicitly estimating the articulation parameters, which improves the smoothness of manipulation. In UMPNet [22], the action is predicted based on two RGBD images: one of the current state and the other of the initial or goal state. Therefore, UMPNet is suitable for goal-conditioned manipulation, aligning closely with the task addressed in our work.
Our approach, however, distinguishes itself from the above methods in three key aspects: (1) While these works rely on single-frame observations to predict actions, our method leverages multi-frame observations through active interaction with the object, enabling more accurate articulation structure estimation. (2) These works use a suction gripper to establish a stable contact with the object and cannot be easily extended to other end-effectors. In contrast, our method supports multiple kinds of end effectors, including suction grippers, two-finger grippers and multi-finger hands. (3) In contrast to articulation flow which represents a one-step object motion in current state, our method builds an explicit model of the articulated object within a physics simulation, which can generate long-horizon multi-step manipulation trajectories prior to physical interaction, and further generalize to unseen interactions, like manipulation with tools demonstrated in our experiments.
II-B Dexterous Manipulation
@@
II-C World Model Construction
@@
In our work, we focus on articulated object manipulation, so we introduce the knowledge prior of the environment by using an explicit physics model. Therefore, we are able to significantly decrease the number of samples required for model construction. Moreover, the explicit physics model’s generalizability guarantees that while we only use a simple action to collect the sample, the built model can be used for long-horizon complex trajectory planning composed of unseen robot actions.
Articulated object modeling has recently gained significant attention in the fields of robotics and computer vision [15, 44, 45, 46]. Ditto [15] employs a 3D network to predict the voxel occupancy and joint parameters simultaneously using two frames of point clouds as input. However, Ditto is limited to objects with a single joint. Liu et al. [47] propose to build kinematic structures from point cloud sequences by jointly optimizing part segmentation, transformation, and joint parameters through an energy minimization framework. But the performance of this approach heavily depends on part segmentation quality, which is not always reliable when the sequence only consists of 2-3 frames. In our work, we address this limitation by designing a novel movable part segmentation module. Our module leverages mesh connectivity, foundation vision models, and proprioception from robotic interactions to achieve more accurate and robust segmentation, leading to better kinematic structure estimation. More recently, Articulate-Anything [48] explore the use of vision-language models to generate programs for constructing Unified Robot Description Format (URDF) files from text, images, and videos. However, their approach relies on mesh retrieval from pre-existing datasets to define object parts, which limits geometric diversity and alignment with real-world observations. In contrast, we leverage advanced 3D AIGC techniques to reconstruct meshes that align precisely with observations. This enables accurate estimation of pose, position, and scale, further supporting precise manipulation in our proposed Sim2Real2 pipeline.
II-D Affordance Learning
@@
[…]
II-E Sim2Real for Robot Learning
@@ III METHOD
The goal of our work is goal-conditioned articulated object manipulation, that is to manipulate articulated objects to specified joint states with various robot-effectors in the real world, including suction grippers, two-finger grippers, and multi-finger dexterous hands. Fig. 2 shows an overview of our framework. It consists of three modules: Interactive Perception (Section III-A), Explicit World Model Construction (Section III-B), Sampling-Based Model Predictive Control (Section III-C).
A single observation of an articulated object cannot provide enough information to reveal its full structure. For example, when humans first look at a kitchen door, it is hard to tell whether it has a rotating hinge or a sliding hinge. However, after the door is moved, humans can use the information from the two observations to infer the type and location of the hinge. Inspired by this, the Interactive Perception module proposes an action to alter the joint state of the articulated object based on learned affordances. This action is then executed on the object in the real world, resulting in a changed state and additional frames of observation.
With the multiple frames of observations, the Explicit World Model Construction module (Section III-B) infers the shape and the kinematic structure of the articulated object to construct a digital model. The digital model can be loaded into a physics simulator for the robot to interact with, forming an explicit world model of the environment.
@@ III-A2 Learn from real-world egocentric demonstrations
Acquiring 3D affordance representations through self-supervised interactions in simulation has shown promise as it does not rely on labeled data. However, certain limitation exists: the success of this method hinges on interactive models in simulation. Unfortunately, the availability of simulated datasets for articulated objects is limited, hindering the generation of training data.
@@
Fig. 3 illustrates how we generate a 3D trajectory for real robot manipulation from 2D affordances. Firstly, we capture an RGB image $\boldsymbol{I}_{0}$ and a depth map $\bm{\mathcal{D}}_{0}$ using the mounted RGBD camera, $\boldsymbol{H}_{0}\in\mathbb{R}^{4\times 4}$ is its relative transformation matrix respect to the robot coordinate system. Secondly, we set a virtual camera, the relative transformation matrix of which is $\boldsymbol{H}\in\mathbb{R}^{4\times 4}$. Since the depth of each pixel in $\boldsymbol{I}_{0}$ is known, we can generate the virtual RGB image $\hat{\boldsymbol{I}}_{0}$ by image wrapping. Thirdly, we use $\boldsymbol{I}_{0}$ and $\hat{\boldsymbol{I}}_{0}$ as the input of the affordance model and generate contact points $\boldsymbol{c}_{0}=(u_{0},v_{0})$, $\hat{\boldsymbol{c}}_{0}=(\hat{u}_{0},\hat{v}_{0})$ and post-contact trajectories $\boldsymbol{\tau}_{0}=(u_{0}^{{}^{\prime}}-u_{0},v_{0}^{{}^{\prime}}-v_{0})$ and $\hat{\boldsymbol{\tau}}_{0}=(\hat{u}_{0}^{{}^{\prime}}-\hat{u}_{0},\hat{v}_{0}^{{}^{\prime}}-\hat{v}_{0})$. The camera intrinsic matrix is $\boldsymbol{K}$, the contact point in the mounted camera frame is $\boldsymbol{p}_{c}\in\mathbb{R}^{3}$, the 3D post contact vector in the camera frame is $\boldsymbol{\tau}_{c}\in\mathbb{R}^{3}$. Fourthly, we respectively calculate the 3D contact point and post contact vector. We use contact point $\boldsymbol{c}_{0}$ to acquire 3D contact point $\boldsymbol{p}_{c}\in\mathbb{R}^{3}$ in the robot base frame:
@@
where $z_{c}$ represents the depth of $\boldsymbol{p}_{c}$. We use the camera’s intrinsic matrix to transfer $\boldsymbol{c}_{0}$ to point in mounted camera frame then use the mounted camera’s extrinsic matrix to transfer it to the 3D point cloud $\boldsymbol{p}_{c}$.
@@
Specifically, our method of calculating 3D post contact vector is shown in Fig. 3. We respectively denote the projection plane of $\boldsymbol{I}_{0}$ and $\hat{\boldsymbol{I}}_{0}$ as $\boldsymbol{S}_{0}$ and $\hat{\boldsymbol{S}}_{0}$. For $\boldsymbol{S}_{0}$, we use $\boldsymbol{\varphi}_{0}$ and $\boldsymbol{\varphi}_{0}^{{}^{\prime}}$ to represent the projection plane. $\boldsymbol{\varphi}_{0}$ represents one possible 3D vector on projection plane $\boldsymbol{S}_{0}$. Its starting point is $\boldsymbol{p}_{c}$, while its ending point can be calculated with:
@@
where $\boldsymbol{o}_{c0}$ is the coordinate of camera frame’s origin in the robot base frame. Then we calculate the norm vector of $\boldsymbol{S}_{0}$: $\boldsymbol{n}_{0}$ = $\boldsymbol{\varphi}_{0}\times\boldsymbol{\varphi}_{0}^{{}^{\prime}}$. We can calculate $\hat{\boldsymbol{n}}_{0}$ in the same way. Finally, we generate the 3D post-contact vector in the robot base frame: $\boldsymbol{\tau}_{c}$ = $\boldsymbol{n}_{0}\times\hat{\boldsymbol{n}}_{0}$.
Finally, we use motion planning to conduct the one-step interaction with the articulated object. The motion planning process can be divided into two phases: we first moves the end effector to the contact point and then we control the end effector to move a certain distance in the direction of the post contact vector. Since the purpose of the one-step interaction is to change the state of the movable part, a pushing interaction is sufficient in most cases. For objects where pushing is not feasible—such as fully closed cabinets and drawers— we take an alternative grasping approach. We first utilize VRB to detect the handle. Next, we open the effector and move it to a position near, but behind, the handle. Then, we move the effector to the handle and close it to grasp the handle. Finally, we move a certain distance in the direction of the post-contact vector. For objects with $K>1$ movable parts, we iteratively detect affordance points by identifying one affordance point, masking out its neighborhood, and then detecting the next affordance point. This process is repeated until $K$ affordance points are identified.
@@ III-B Explicit World Model Construction
Figure 4: Our pipeline for explicit world model construction: For each state of the articulated object, we begin by generating an unaligned and unscaled mesh from multi-view RGB images using 3D AIGC. Next, we estimate the scale and pose through differentiable rendering, and segment the aligned mesh into sub-parts. Once segmented point clouds for each state are obtained, we infer movable part segmentation by analyzing differences between frames of point clouds. We estimate the kinematic structure of the mesh, including the part tree hierarchy, joint categories (prismatic or revolute), and joint configurations (axis direction and origin). Finally, we construct a digital twin of the articulated object represented in URDF format, which can be easily loaded into different physics simulators.
In this section, we propose a novel explicit world model construction method that we reconstruct the geometry and the kinematic structure of an articulated object. Fig. 4 shows the whole pipeline. We assume the object consists of one root part and $K$ movable parts. The input are $K+1$ frames of observations:$\boldsymbol{I}^{(k)}_{0}$, $\boldsymbol{D}^{(k)}_{0}$, $\boldsymbol{M}^{(k)}_{0}$, $\boldsymbol{I}^{(k)}_{1}$, $\boldsymbol{M}^{(k)}_{1}$,…, $\boldsymbol{I}^{(k)}_{C}$, $\boldsymbol{M}^{(k)}_{C}$}, $k=0,...,K$. $\boldsymbol{I}^{(k)}_{0}$, $\boldsymbol{D}^{(k)}_{0}$ are the RGB image and depth map from the main RGBD sensor, and $\boldsymbol{M}^{(k)}_{0}$ is the object segmentation mask of $\boldsymbol{I}^{(k)}_{0}$. Besides the main sensor, we also use $C$ RGB cameras located at different viewpoints to improve the reconstruction quality. $\boldsymbol{I}^{(k)}_{c}$ is the RGB image captured by the $c$th camera, and $\boldsymbol{M}^{(k)}_{c}$ is the corresponding object segmentation mask. In this work, we use $C=1$.
III-B1 Canonical Mesh Generation
The first step of the pipeline is to reconstruct the $k$th frame canonical mesh $\bm{\widetilde{\mathcal{A}}}^{(k)}$. Here we utilize a controllable large-scale generative model, Rodin [64], to generate $\bm{\widetilde{\mathcal{A}}}^{(k)}$ conditioned on $\boldsymbol{I}^{(k)}_{0}$, $\boldsymbol{M}^{(k)}_{0}$, $\boldsymbol{I}^{(k)}_{1}$, $\boldsymbol{M}^{(k)}_{1}$,…, $\boldsymbol{I}^{(k)}_{C}$, $\boldsymbol{M}^{(k)}_{C}$}.
III-B2 Mesh Positioning and Scaling
While the generated mesh $\bm{\widetilde{\mathcal{A}}}^{(k)}$ exhibits high-quality geometry and texture, it is positioned at the origin and scaled to a unit size. Therefore, it is required to obtain the real scale and position of the object for further operation.
To achieve this, we first estimate the position of the object. Since $\bm{\widetilde{\mathcal{A}}}^{(k)}$ is scaled to unit size, we cannot directly apply CAD-model-based 6D pose estimation methods like FoundationPose [65]. Therefore, we modify FoundationPose to meet our specific requirements. Specifically, we first render a set of images of $\bm{\widetilde{\mathcal{A}}}^{(k)}$ with different pose hypotheses and sort these poses based on the images’ similarity to $\boldsymbol{I}^{(k)}_{0}$. Secondly, for the pose with the highest similarity, we refine it by aligning the rendered edges with the edges of $\boldsymbol{M}^{(k)}_{0}$ using differentiable rendering [66]. To improve the stability of the optimization, we constrain the orientation to the $Z$-axis.
After determining the orientation $\bm{R}$ and the unscaled translation $\bm{\widetilde{t}}$ , we can render an unscaled depth map $\widetilde{\boldsymbol{D}}^{(k)}_{0}$. Then the scale $s^{(k)}$ can be computed as:
$s^{(k)}=\frac{\sum_{\boldsymbol{M}^{(k)}_{0}}\boldsymbol{D}^{(k)}_{0}}{\sum_{\boldsymbol{M}^{(k)}_{0}}\widetilde{\boldsymbol{D}}^{(k)}_{0}}$ (5)
Once $s^{(k)}$ is computed, we can generate a correctly scaled mesh $\bm{{\mathcal{A}}}^{(k)}$ and position it accurately to align with its real-world location.
III-B3 Movable Part Segmentation
To build an interactable model of the object, we need to segment the mesh $\bm{{\mathcal{A}}}^{(k)}$ into movable parts. However, it is difficult to distinguish the parts from a single-frame static mesh. Therefore, we use all the $K+1$ meshes with different kinematic states to segment the parts. The output is a point cloud $\bm{{\mathcal{P}}}^{(K)}\in\mathbb{R}^{N^{(K)}\times 3}$ sampled from $\bm{{\mathcal{A}}}^{(K)}$ with part segmentation labels $\bm{{\mathcal{S}}}^{(K)}\in\{0,1,...,K\}^{N^{(K)}}$, where $N^{(K)}$ is the number of points of $\bm{{\mathcal{P}}}^{(K)}$.
We first segment the mesh into small sub-parts $\bm{{\mathcal{Q}}}^{(k)}$ using foundation vision models. In this work, we choose SAMPart3D [67]. To determine whether each sub-part is moved, we find the sub-part $\bm{P}$ ’s closest point set $\bm{PB}$ in $\bm{{\mathcal{P}}}^{(k-1)}$ and compute the chamfer distance $\{d_{i}^{(k)\rightarrow(k-1)}\}$, then we find $\bm{PB}$’s closest points $\bm{PF}$ in $\bm{{\mathcal{P}}}^{(k)}$ and compute the distances $\{d_{i}^{(k-1)\rightarrow(k)}\}$. We set $\bm{P}$ as moved if $\{d_{i}^{(k)\rightarrow(k-1)}\}$ is larger than certain threshold or the distribution of $\{d_{i}^{(k)\rightarrow(k-1)}\}$ and $\{d_{i}^{(k-1)\rightarrow(k)}\}$ are different. However, this criteria may result in wrong segmentation due to the inconsistency of 3D AIGC for different frames or the error of pose estimation.
[…]
Algorithm 1 Movable Part Segmentation
0:  List of meshes $\{\bm{{\mathcal{A}}}^{(k)}\}_{k=0}^{K}$
0:  Movable part segmentation $\bm{{\mathcal{S}}}^{(K)}$
1:  for $k=0,\dots,K$ do
2:     Segment $\bm{{\mathcal{A}}}^{(k)}$ into small sub-parts $\bm{{\mathcal{Q}}}^{(k)}$ using SAMPart3D [67]
[…]
4:     Sample point cloud $\bm{{\mathcal{P}}}^{(k)}$ with sub-parts indices from $\bm{{\mathcal{A}}}^{(k)}$
5:  end for
6:  $\bm{{\mathcal{S}}}^{(0)}=0$
7:  for $k=1,\dots,K$ do
8:     # Breadth-first traversal for graph $\bm{{\mathcal{G}}}^{(k)}$ starting from $\bm{Q}_{\text{closest}}$
9:     for Sub-part $\bm{Q}$ do
10:        # Determine whether each sub-part is moved
11:        Get point indices ${\bm{x}}$
12:        Sub-part point cloud $\bm{P}=\bm{{\mathcal{P}}}^{(k)}[\bm{x}]$
13:        Find $\bm{P}$’s closest points $\bm{PB}$ in $\bm{{\mathcal{P}}}^{(k-1)}$ and compute the distances $\{d_{i}^{(k)\rightarrow(k-1)}\}$
14:        Find $\bm{PB}$’s closest points $\bm{PF}$ in $\bm{{\mathcal{P}}}^{(k)}$ and compute the distances $\{d_{i}^{(k-1)\rightarrow(k)}\}$
15:        Compute whether $\bm{Q}$ is moved using $\{d_{i}^{(k)\rightarrow(k-1)}\}$ and $\{d_{i}^{(k-1)\rightarrow(k)}\}$
16:        Compute if at least one predecessor sub-part of $\bm{Q}$ is moved
17:        if $\bm{Q}$ is moved then
18:           $\bm{{\mathcal{S}}}^{(k)}[\bm{x}]=k$
19:        else
20:           $\bm{{\mathcal{S}}}^{(k)}[\bm{x}]=\bm{{\mathcal{S}}}^{(k-1)}[\bm{x}_{\text{closest}}]$
21:        end if
22:     end for
23:  end for
24:  return  $\bm{{\mathcal{S}}}^{(K)}$
III-B4 Kinematic Structure Construction
The final procedure is to construct the object’s kinematic structure, including the part tree structure, joint categories (prismatic or revolute), joint configurations (axis direction and origin). Here we adopt  [47] to estimate the kinematic structure, where we first estimate the per-part rigid body transformations and further project them into a valid kinematic model.